A 435−g piece of copper tubing is heated to 89.5°C and placed in an insulated vessel containing 159 g of water at 22.8°C. Assuming no loss of water and a heat capacity for the vessel of 10.0 J/°C, what is the final temperature of the system (c of copper = 0.387 J/g·°C)?

1
Expert's answer
2021-03-08T06:11:05-0500

-q(Cu) = q(H2O) + q(vessel)

q = m * c * (T2 - T1)

-435 * 0.387 * (T - 89.5) = 159 * 4.2 * (T - 22.8) + 10.0 * (T - 22.8)

-168.345*T +15066.8775 = 667.8*T - 15225.84 + 10*T - 228

30520.7175 = 846.145*T

T = 36.07°C ---- answer